OpenStudy (anonymous):

What is the x-intercept for y = (x – 1)^2 – 16? I already know the y-intercept is -16.

OpenStudy (anonymous):

set equal to zero and solve \[(x-1)^2-16=0\] \[(x-1)^2=16\] \[x-1=4\iff x=5\] or \[x-1=-4\iff x=-3\] \

OpenStudy (anonymous):

\[(x-1)^2 - 16\] \[(x-1)^2 = 16\] \[x^2-2x+1 = 16\] \[x^2-2x = 16-1\] \[x^2 - 2x = 15\] \[x(x-2)=15\] so..\[x= 15 \] and \[x-2 = 15\] \[x=15+2= 17\] x=15 and x=17